I am having a hard time figuring out whether to add or subtract this system of equations (must solve by elimination).
5x+3y=8
2x-4y=-6
I multiplied the first equation by 4 and the second by 3 and I ended up with 20x+12y=32
6x-12y=-18
26x=14
11x=7
x=7/11 I don't think this is right and I can't solve for y. Any help would be greatly appreciated.
Thanks

Answer by stanbon(75887)

I am having a hard time figuring out whether to add or subtract this system of equations (must solve by elimination).
5x+3y=8
2x-4y=-6
I multiplied the first equation by 4 and the second by 3 and I ended up with 20x+12y=32
6x-12y=-18
26x=14
x = 7/13
-----
Solve for "y":
5x + 3y = 8
5(7/13) + 3y = 8
(35/13) + 3y = (104/13)
3y = (69/13)
y = 23/13
==================
Cheers,
Stan H.